# The sPHENIX Micromegas Outer Tracker

### v1

@@ 2.5.3 The 256-channel Front-End Electronics (FEE) Board
Each FEE board has eight SAMPA v5 chips that accept 256 inputs from the detector. The block diagram of the FEE board is shown in Figure 11. A picture of the board is shown in Figure 12. The two SAMTEC SEAF connectors on the right collect the signal. They connect to the Micromegas chamber strips using a dedicated transition cable, shown in blue in Figure 8. Connectors on the right, top-to-bottom, correspond to the optical transceiver, low voltage input, and a JTAG interface for board configuration redundant with the interface implemented over the optical link. A Xilinx Artix-7 200T FPGA (XC7A200T) mounted on the FEE has multiple functions: it collects data from the SAMPA chips, formats them into a data packet, and sends the packet off to the back-end electronics (DAM/EBDC) via the optical transceiver.
@@ 2.5.4 Magnetic field and radiation tolerance
Radiation tolerance for the FEE is another key factor for validating its design and selection of individual electronic parts. The same FEE is used for TPOT and the TPC. For the TPC, some boards are installed close to the beam pipe, where the radiation dose is maximum. The Total Ionization Dose (TID) at a 16 cm distance from the beam pipe and for 5 years of sPHENIX data taking is estimated to be 25 kRad, using data from the PHENIX experiment ADCOX2003469 . It is significantly smaller at the TPOT location. TID tests performed on the optical transceiver and the full FEE board using a ${}^{60}$o $\gamma$ source show that all components of the FEE can sustain a 100 kRad dose except the EEPROM (42 kRad) and the Phase-Lock Loop (PLL, $\sim 50$ kRad). These limits are well above the dose expected during sPHENIX running.
@@ 2.5.5 Mighty-Jack and Sort-out-Box
Figure 13: Left: Picture of the Mighty-JACK board, right: picture of the Sort-Out-Box (SOB).
@@ 2.5.6 The back-end electronics (DAM and EBDC)
The FELIX board was developed primarily for the ATLAS experiment at CERN Ryu_2017 , with the help of the BNL Instrumentation Division and the ATLAS group at BNL. It satisfies all the requirements for interfacing with the SAMPA FEE. A picture of the DAM board is shown in Figure 15, and Figure 16 shows the corresponding block diagram, both using ATLAS FELIX v2.1 card.
@@ 2.6.3 Cooling
Figure 18: Left: design of the FEE cooling plates; right: picture of one assembled cooling plate.
The FEE boards are cooled using a sub-atmospheric pressure cold water circuit. The water is circulated through aluminum cooling plates mounted on top of each board and separated by a thermally conductive gap-pad material. A diagram of the cooling plate and a picture of the assembled plate are shown in Figure 18. A Childyne unit is used to circulate the sub-atmospheric pressure water. Water is provided to each module independently and each line is monitored by its individual flowmeter. The cooling plates of the two FEE boards of a given module are mounted in series. Using sub-atmospheric pressure water ensures that in the event of a leak, air leaks inward but no liquid leaks outward. With this setup, the cooling water input temperature set to 21°$\degree$°C and the cooling water flow rate to 125c̃cm, the temperature measured on the FPGA of the FEE boards is around 40°$\degree$°C.
@@ 2.6.4 Gas system
Figure 19 shows the schematics of the TPOT gas system. The two chambers in each module are connected in series. Each module has a volume of $1.4$ l for a total of $11.2$ l. The gas flow is operated in the 50-100 ccm range. Each module is equipped with its own flowmeter.
@@ 2.8 Slow Control and monitoring
The TPOT slow control and monitoring system tracks a number of measurements relevant to detector operation using a Grafana web interface and the time series database Prometheus Web:Grafana:Docs . This includes, but is not limited to: (i) the voltage and current of each HV channel, (ii) the voltage and current of each LV channel, (iii) the log and history of tripped HV channels, (iv) the temperature of the FEE and (v) the recorded data rate. Five temperatures per FEE are measured to monitor the FPGA as well as the FEE PCB. The LV power supply is interlocked to the sub-atmospheric cooling water system, so that it is not possible to turn ON the FEE in absence of cooling. An example of the monitoring of HV and current values over time is shown in Figure 21.
@@ 3.2 Metrology and Survey
The combination of these three sets of measurements allows the position of the strips inside sPHENIX to be known at a precision of a few 100 $\mu$m. This determination is then improved using the reconstructed trajectory of particles passing through sPHENIX tracking detectors, to an accuracy of 100 $\mu$m or less. Millepede Blobel:2002ax is used to perform this track-based alignment.
@@ 4.1 Detector characterization at Saclay
Figure 24 shows a picture of the cosmic test bench used at Saclay for the characterization of the TPOT modules during production. The modules are positioned in the middle trays of the setup between two sets of MultiGen detectors  BOUTEILLE2016187 (two detectors above and two detectors below), used for the reconstruction of the reference track. 3D-printed indentation fixtures are attached to the cosmic test bench so all the modules are tested at the same position with respect to the reference detectors. Data acquisition is triggered using a coincidence of scintillating pads mounted on Photo-Multiplier Tubes (PMT). The signal is read with the DREAM electronics flouzat2014dream originally developed for the Micromegas detector of the CLAS12 experiment at the Jefferson Laboratory BURKERT2020163419 ; osti_1607856 .
@@
Figure 25: Detection efficiency as a function of position in two chambers of the same TPOT module. Left: $\phi$ view; right: $z$ view.
@@
Figure 27: Distribution of the residuals (cluster-reference) in two chambers of the same TPOT module. Left: $\phi$ view; right: $z$ view.
Figure 27 shows the residuals measured in the two chambers of a given TPOT module, obtained by comparing the position of the measured cluster to that of the track provided by the reference detectors of the cosmic test bench. The width of the residuals distribution is the quadratic sum of the chamber intrinsic spatial resolution and the accuracy of the reference track. It is an upper limit to the chamber intrinsic resolution. The sum of two Gaussian distributions is used to fit the signal. The width of the narrow Gaussian distribution is 390 $\mu$m (500 $\mu$m) for the $\phi$ ($z$) view, respectively. The larger value measured for the $z$ view is due to its larger pitch. The wide Gaussian distribution is used to describe the tails of the residuals distribution. It has a width of about 1.5 mm for both views. These tails are attributed in parts to (i) the contamination from poorly defined reference cosmic tracks, (ii) the contamination from poorly defined clusters in the detector, and (iii) the presence of the pillars that support the Micromegas mesh at a fixed distance above the resistive layer and distort the electric field locally.
@@ 4.2 Detector characterization at BNL
4.2.1 Detection efficiency with SAMPA electronics
The detector performance presented in section 4.1 was measured using the data acquisition system available at Saclay, and in particular the DREAM ASIC flouzat2014dream . Compared to SAMPA, primarily designed for low-capacitance detectors (a few 10 pF), DREAM is better suited to the typical capacitance of Micromegas detectors (up to 300 pF in the TPOT case) leading to improved signal-to-noise ratio, smaller signal detection threshold and a higher detection efficiency at a given amplification gain.
@@ 4.3.2 Operation Point Determination
With magnetic field: the sPHENIX magnetic field is perpendicular to the drift direction of the primary electrons deposited by particles in the Micromegas drift space. Consequently it curves the trajectory of the primary electrons in the transverse plane, resulting in a non-zero value for what is known as the Lorentz angle, which is the angle between the drift direction of the electrons and the direction of the electric field. Too large values of the Lorentz angle result in deteriorated spatial resolution and efficiency due to spread of primary electrons. For a given magnetic field, the Lorentz angle is reduced by increasing the drift electric field. This is achieved by increasing the drift HV.
@@
Figure 32 shows an estimate of the detection efficiency measured in the two chambers of the same TPOT module as a function of the resistive layer HV with magnetic field on and a drift electrode HV of -400 V. The boxes around the points represent the systematic uncertainties on the measurement, estimated by varying the selection criteria of either the reference or the measured cluster sample. They are largely correlated point-to-point. Full efficiency is reached starting from approximately 410 V. The detection efficiency dependence on the resistive layer voltage is similar to that shown in Figure 26 (right) and measured with cosmic tracks. However, i) the efficiency measured on the plateau is significantly lower (about $90$% as opposed to $100$% in Fig. 32 and ii) the efficiency does not drop to zero for small values of the voltage. Both observations are attributed to the poor definition of the reference sample. A more accurate determination of the reference sample and the detection efficiency requires correlating the TPOT measurements to that of other tracking subsystems, in particular the TPC.
@@ 4.3.3 Operation Mode Description
Figure 33 shows the cluster size distribution and the number of signal hits per ${\rm Au+Au}$ collision in TPOT. The mean cluster size is 3.5 (2.6) for $\phi$ ($z$) views. The larger size measured in the $\phi$ views is attributed to the detector pitch being smaller and the azimuthal angular distribution of the particles passing through the chamber being wider due to the presence of the magnetic field. The mean occupancy, defined as the mean fraction of strips with a signal above threshold in a given chamber and per ${\rm Au+Au}$ collision, is $9.3\%$ ($6.2$) for $\phi$ ($z$) views. The difference between the two is attributed to the larger mean cluster size measured for $\phi$ views.
@@ 4.3.4 Correlation with Other Subsystems
Three examples of cosmic tracks reconstructed in MVTX, INTT and TPOT are shown in Figure 36. The curvature of the trajectory is due to the presence of the magnetic field. Successfully reconstructing these trajectories with the MVTX, INTT and TPOT and extrapolating it inside the TPC is key to measuring and correcting for the beam-induced distortions inside the TPC.

### v3

@@ 2.5.3 The 256-channel Front-End Electronics (FEE) Board
Each FEE board has eight SAMPA v5 chips that accept 256 inputs from the detector. The block diagram of the FEE board is shown in Figure 11. A picture of the board is shown in Figure 12. The two SAMTEC SEAF connectors on the left collect the signal. They connect to the Micromegas chamber strips using a dedicated transition cable, shown in blue in Figure 4. Connectors on the right side of the board in Figure 12, top-to-bottom, correspond to the optical transceiver, low voltage input, and a JTAG interface for board configuration redundant with the interface implemented over the optical link. A Xilinx Artix-7 200T FPGA (XC7A200T) mounted on the FEE has multiple functions: it collects data from the SAMPA chips, formats them into a data packet, and sends the packet off to the back-end electronics (DAM/EBDC) via the optical transceiver.
@@ 2.5.4 Magnetic field and radiation tolerance
Radiation tolerance for the FEE is another key factor for validating its design and selection of individual electronic parts. The same FEE is used for TPOT and the TPC. For the TPC, some boards are installed close to the beam pipe, where the radiation dose is maximum. The Total Ionization Dose (TID) at a 16 cm distance from the beam pipe and for 5 years of sPHENIX data taking is estimated to be 25 kRad, using data from the PHENIX experiment [12]. It is significantly smaller at the TPOT location. TID tests performed on the optical transceiver and the full FEE board using a 60Co $\gamma$ source show that all components of the FEE can sustain a 100 kRad dose except the EEPROM (42 kRad) and the Phase-Lock Loop (PLL, $\sim 50$ kRad). These limits are well above the dose expected during sPHENIX running.
@@ 2.5.5 Mighty-Jack and Sort-out-Box
Figure 13: Picture of the Mighty-JACK board integrated into the Sort-Out-Box (SOB).
@@ 2.5.6 The back-end electronics (DAM and EBDC)
The FELIX board was developed primarily for the ATLAS experiment at CERN [7], with the help of the BNL Instrumentation Division and the ATLAS group at BNL. It satisfies all the requirements for interfacing with the SAMPA FEE. A picture of the DAM board is shown in Figure 15, and Figure 16 shows the corresponding block diagram, both using ATLAS FELIX v2.1 card.
@@ 2.6.3 Cooling
Figure 18: Left: design of the FEE cooling plates. Shown, right-to-left, are: the top aluminum plate, the aluminum tube through which sub-atmospheric cooling water is circulated, a set of washers used to control the thickness of the glue used to assemble the cooling plate (832TC-B thermally conductive epoxy, from MG Chemicals), the bottom aluminum plate, and the thermally conductive 2.5 mm thick gap-pad material layer (Tflex 300 Series from Laird Technologies); Right: picture of one assembled cooling plate.
The FEE boards are cooled using a sub-atmospheric pressure cold water circuit. The water is circulated through aluminum cooling plates mounted on top of each board and separated by a thermally conductive gap-pad material of thickness 2.5 mm. A diagram of the cooling plate and a picture of the assembled plate are shown in Figure 18. The gap-pad material is from Laird Technologies (Tflex 300 Series). It is made of silicone elastomer, has a density of 1.78 g/cc and a thermal conductivity of 1.2 W/mK. A Childyne unit is used to circulate the sub-atmospheric pressure water. Water is provided to each module independently and each line is monitored by its individual flowmeter. The cooling plates of the two FEE boards of a given module are mounted in series. Using sub-atmospheric pressure water ensures that in the event of a leak, air leaks inward but no liquid leaks outward. With this setup, the cooling water input temperature set to 21°$\degree$°C and the cooling water flow rate to 125 cc/min, the temperature measured on the FPGA of the FEE boards is around 40°$\degree$°C.
@@ 2.6.4 Gas system
Figure 19 shows the schematics of the TPOT gas system. The two chambers in each module are connected in series. Each module has a volume of $1.4$ l for a total of $11.2$ l. The gas flow is operated in the 50-100 cc/min range. Each module is equipped with its own flowmeter.
@@ 2.8 Slow Control and monitoring
The TPOT slow control and monitoring system tracks a number of measurements relevant to detector operation using a Grafana web interface and the time series database Prometheus [13]. This includes, but is not limited to: (i) the voltage and current of each HV channel, (ii) the voltage and current of each LV channel, (iii) the log and history of tripped HV channels, (iv) the temperature of the FEE and (v) the recorded data rate. Five temperatures per FEE are measured to monitor the FPGA as well as the FEE PCB. The LV power supply is interlocked to the sub-atmospheric cooling water system, so that it is not possible to turn ON the FEE in absence of cooling. An example of the monitoring of HV and current values over time is shown in Figure 21.
@@ 3.2 Metrology and Survey
The combination of these three sets of measurements allows the position of the strips inside sPHENIX to be known at a precision of a few 100 $\mu$m. This determination is then improved using the reconstructed trajectory of particles passing through sPHENIX tracking detectors, to an accuracy of 100 $\mu$m or less. Millepede [14] is used to perform this track-based alignment.
@@ 4.1 Detector characterization at Saclay
Figure 24 shows a picture of the cosmic test bench used at Saclay for the characterization of the TPOT modules during production. The modules are positioned in the middle trays of the setup between two sets of MultiGen detectors  [15] (two detectors above and two detectors below), used for the reconstruction of the reference track. 3D-printed indentation fixtures are attached to the cosmic test bench so all the modules are tested at the same position with respect to the reference detectors. Data acquisition is triggered using a coincidence of scintillating pads mounted on Photo-Multiplier Tubes (PMT). The signal is read with the DREAM electronics [16] originally developed for the Micromegas detector of the CLAS12 experiment at the Jefferson Laboratory [17, 18].
Figure 25: Detection efficiency as a function of position in two chambers of the same TPOT module, for a resistive layer HV of 410 V. Left: $\phi$ view; right: $z$ view.
@@
Figure 27: Distribution of the residuals (cluster-reference) in two chambers of the same TPOT module, for a resistive layer HV of 410 V. Left: $\phi$ view; right: $z$ view.
Figure 27 shows the residuals measured in the two chambers of a given TPOT module, obtained by comparing the position of the measured cluster to that of the track provided by the reference detectors of the cosmic test bench. The width of the residuals distribution is the quadratic sum of the chamber intrinsic spatial resolution, the accuracy of the reference track and the contribution from multiple scattering in the detectors. The combined contribution from multiple scattering and track accuracy is estimated to be 200 $\mu$m (about 150 $\mu$m each). The sum of two Gaussian distributions is used to fit the signal. The width of the narrow Gaussian distribution is 390 $\mu$m (420 $\mu$m) for the $\phi$ ($z$) view, respectively, corresponding to estimated detector resolutions of 330 $\mu$m (370 $\mu$m). The larger value measured for the $z$ view is due to its larger pitch. The wide Gaussian distribution is used to describe the tails of the residuals distribution. It has a width of about 1.2 mm for both views.
All the Micromegas chambers build at Saclay for the TPOT detector have been characterized on the cosmic test bench shown in Figure 24, using the same, systematic procedure. Performances similar to those shown in Figures 25, 26 (left panel) and 27 have been measured for the 16 chambers used in the TPOT detector. Table 1 shows the detection efficiencies and leak rates measured for nine out of the ten TPOT modules constructed at Saclay. The first module, Module 1, not listed in the table, was sent early on to BNL for testing with the SAMPA electronics (Section 4.2). It was therefore not characterized in the same manner, nor used for the final TPOT detector. Module 2, which shows a slightly lower efficiency for the $z$ view was not used for the TPOT detector either.
Module Resist HV Eff. ($\phi$) Eff. ($z$) Max. leak rate
(V) (%) (%) (l/h)
Module 2 420 99.2 92.9 0.03
Module 3 420 98.7 98.1 0.03
Module 4 430 99.1 98.3 0.02
Module 5 430 99.0 98.1 0.03
Module 6 420 98.2 98.2 0.03
Module 7 420 99.6 97.9 0.01
Module 8 430 99.1 98.6 0.03
Module 9 430 99.1 98.7 0.04
Module 10 430 99.1 98.8 0.04
Table 1: Detection efficiency and leak rate measured at Saclay for nine out of ten TPOT modules, during detector fabrication.
@@ 4.2 Detector characterization at BNL
The detector performance presented in section 4.1 was measured using the data acquisition system available at Saclay, and in particular the DREAM ASIC [16]. Compared to SAMPA, primarily designed for low-capacitance detectors (a few 10 pF), DREAM is better suited to the typical capacitance of Micromegas detectors (up to 300 pF in the TPOT case) leading to improved signal-to-noise ratio, smaller signal detection threshold and a higher detection efficiency at a given amplification gain.
@@ 4.3.2 Operation Point Determination
With magnetic field: the sPHENIX magnetic field is perpendicular to the drift direction of the primary electrons deposited by particles in the Micromegas drift space. Consequently it curves the trajectory of the primary electrons in the transverse plane, resulting in a non-zero Lorentz angle, defined as the angle between the drift direction of the electrons and the direction of the electric field. Too large values of the Lorentz angle result in deteriorated spatial resolution and efficiency due to spread of primary electrons. For a given magnetic field, the Lorentz angle is reduced by increasing the drift electric field. This is achieved by increasing the drift HV.
@@
Figure 32 shows an estimate of the detection efficiency measured in the two chambers of the same TPOT module as a function of the resistive layer HV with magnetic field on and a drift electrode HV of -400 V. The boxes around the points represent the systematic uncertainties on the measurement, estimated by varying the selection criteria of either the reference or the measured cluster sample. They correspond to the best accuracy with which the detection efficiency of the TPOT chambers could be measured in ${\rm Au+Au}$ collisions and during data taking. They are largely correlated point-to-point. Full efficiency is reached starting from approximately 410 V. The detection efficiency dependence on the resistive layer voltage is similar to that shown in Figure 26 (right) and measured with cosmic tracks. However, i) the efficiency measured on the plateau is lower (about 90$\pm$7as opposed to $100$% in Figure 32) and ii) the efficiency does not drop to zero for small values of the voltage. Both observations are attributed to the poor definition of the reference sample. A more accurate determination of the reference sample and the detection efficiency requires correlating the TPOT measurements to that of other tracking subsystems, in particular the TPC.
@@ 4.3.3 Operation Mode Description
Figure 33 shows the cluster size distribution and the number of signal hits per ${\rm Au+Au}$ collision in TPOT. The mean cluster size is 3.5 (2.6) for $\phi$ ($z$) views. The larger size measured in the $\phi$ views is attributed to the detector pitch being smaller and the azimuthal angular distribution of the particles passing through the chamber being wider due to the presence of the magnetic field. Similar cluster sizes, within 5%, are measured for all chambers of the same strip orientation. They are larger than those measured with cosmic data (1.8 and 1.6 for $\phi$ and $z$ views, respectively). This increase is attributed to the angular distribution of particles passing through the detector as well as their mean deposited energy being larger in ${\rm Au+Au}$ collisions than with cosmics. The mean occupancy, defined as the mean fraction of strips with a signal above threshold in a given chamber and per ${\rm Au+Au}$ collision, is $9.3\%$ ($6.2$) for $\phi$ ($z$) views. The difference between the two is attributed to the larger mean cluster size measured for $\phi$ views.
@@ 4.3.4 Correlation with Other Subsystems
Three examples of cosmic tracks reconstructed in MVTX, INTT and TPOT are shown in Figure 36. The curvature of the trajectory is due to the presence of the magnetic field. Successfully reconstructing these trajectories with the MVTX, INTT and TPOT and extrapolating it inside the TPC is key to measuring and correcting for the beam-induced distortions of the electron drift inside the TPC.